OpenStudy (cutiecomittee123):

A bowl of soup cools according to Newtons law of cooling. Its temperature (in degrees farenheit) at time t is given by T(t)=68+144e^0.04t. Where t is given in minutes. How long after serving is the soup 125 degrees?

jimthompson5910 (jim_thompson5910):

it should be \[\Large T(t)=68+144e^{-0.04t}\] the exponent is -0.04t and not 0.04t

jimthompson5910 (jim_thompson5910):

Replace T(t) with 125 and solve for t \[\Large T(t)=68+144e^{-0.04t}\] \[\Large 125=68+144e^{-0.04t}\] \[\Large 125-68=144e^{-0.04t}\] \[\Large 57=144e^{-0.04t}\] \[\Large \frac{57}{144} = e^{-0.04t}\] \[\Large 0.3958333 = e^{-0.04t}\] I'll let you finish up

OpenStudy (cutiecomittee123):

Thank you.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (cutiecomittee123):

wait so I get .395/-0.04 which is -9 something, but you cant have negatives as minutes that have passed

OpenStudy (cutiecomittee123):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how do you undo exponents? which type of function do you need to apply?

OpenStudy (cutiecomittee123):

I tried rounding uo to .4/-0.04 but I just got -10

OpenStudy (cutiecomittee123):

Oh yeah you have to use logs

jimthompson5910 (jim_thompson5910):

specifically, log with base e aka LN \[\Large \log_{e} = \ln\]

OpenStudy (cutiecomittee123):

So then ln(.395)/-0.04?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

what is that equal to?

OpenStudy (cutiecomittee123):

23.22 minutes:)

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (cutiecomittee123):

sweet.